# ToddlerBot: Open-Source ML-Compatible Humanoid Platform for Loco-Manipulation

### v2

@@ Abstract
Learning-based robotics research driven by data demands a new approach to robot hardware design—one that serves as both a platform for policy execution and a tool for embodied data collection to train policies. We introduce ToddlerBot, a low-cost, open-source humanoid robot platform designed for scalable policy learning and research in robotics and AI. ToddlerBot enables seamless acquisition of high-quality simulation and real-world data. The plug-and-play zero-point calibration and transferable motor system identification ensure a high-fidelity digital twin, enabling zero-shot policy transfer from simulation to the real-world. A user-friendly teleoperation interface facilitates streamlined real-world data collection for learning motor skills from human demonstrations. Utilizing its data collection ability and anthropomorphic design, ToddlerBot is an ideal platform to perform whole-body loco-manipulation. Additionally, ToddlerBot’s compact size ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensures safe operation in real-world environments. Reproducibility is achieved with an entirely 3D-printed, open-source design and commercially available components, keeping the total cost under $6000\leavevmode\nobreak\ \mathrm{USD}$. Comprehensive documentation allows assembly and maintenance with basic technical expertise, as validated by a successful independent replication of the system. We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, loco-manipulation tasks, and a collaborative long-horizon scenario where two robots tidy a toy session together. By advancing ML-compatibility, capability, and reproducibility, ToddlerBot provides a robust platform for scalable learning and dynamic policy execution in robotics research.
@@
Figure 1: ToddlerBot is an open-source, ML-compatible humanoid platform for efficiently collecting large-scale, high-quality training data in both simulation and the real world. We leverage massive parallel environments, an accurate digital twin for simulation, and an intuitive teleoperation device for precise upper and lower body control in real-world data collection. We demonstrate that ToddlerBot meets these requirements and successfully acquires a diverse set of loco-manipulation skills from both data sources, including walking, push-ups, pull-ups, wagon pushing, bimanual, and full-body manipulation.
I Introduction
@@
Humanoid Size(m) Weight(kg) # Active DoFs(a) Manipulation Locomotion Sim Data Real Data Open Source Price(\$)
BD Atlas [8] 1.50 89.0 28 - - -
Berkeley [27] 0.85 16.0 12 code 10K
Booster T1 [35] 1.18 30.0 23 code 34K
BRUCE [28] 0.70 4.8 16 code 6.5K
Cassie [33] 1.15 35.0 10 code 250K
Digit [34] 1.75 65.0 16 code 250K
Duke [49] 1.00 30.0 10 everything 16K
Figure [9] 1.68 70.0 26 - - -
Fourier GR1 [10] 1.65 55.0 32 - - 110K
iCub [32] 1.04 24.0 32 code 300K
MIT [5] 1.04 24.0 18 - - -
NAO H25 [31] 0.57 5.2 23 code 14K
Optimus [44] 1.73 57.0 28 - - -
Robotis OP3 [36] 0.51 3.5 20 code 11K
Unitree G1 [47] 1.32 35.0 29 code 57K
Unitree H1 [46] 1.76 47.0 19 code 70K
Ours 0.56 3.4 30 everything(b) 6K
Average Adult [14] 1.73 70.9 32(c) - - - -
@@
  • (b)
    Everything includes the digital twin, learning algorithms, hardware design, assembly manual, detailed documentation, and tutorials.
@@
    While human body is powered by over 600 muscles in reality, the primary functional movements of the human body can be approximated using 32 revolute joints: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes.
Table I: Comparison with Other Humanoid Research Platforms.
Conventional robot design prioritizes factors such as actuator strength, sensor accuracy, mechanical precision, and repeatability—key objectives for developing robust control algorithms. However, these platforms are not inherently aligned with modern robot learning paradigms driven by embodied data. A robot platform compatible with a machine learning (ML) approach must possess the innate ability to collect observation and action data seamlessly, both in simulation and in the real world, as these complementary data sources are essential for scalable policy learning. Simulation enables rapid and scalable data collection but relies heavily on accurate physics models. In contrast, real-world data tend to be more reliable but are often difficult to scale due to cost and safety concerns.
While some recent quadrupeds [23, 24] and robotic manipulators [50, 48, 41, 38, 2] have been designed with ML-compatibility in mind, we introduce ToddlerBot, a humanoid robot platform for robotics and AI research community (Figure 1), specifically developed to facilitate policy learning for both locomotion and manipulation skills. ToddlerBot is designed to maximize ML-compatibility, capability, and reproducibility, while minimizing the expertise and costs required for its construction and maintenance.
To address the need for ML-compatibility, ToddlerBot is designed not only to execute policies but also to serve as a robust data collection platform. It enables the acquisition of high-quality simulation data through a plug-and-play zero-point calibration procedure and transferable motor system identification (sysID) results. These tools ensure a high-fidelity digital twin without the need for additional tuning. We validate the quality of simulation data with both keyframe-interpolated motions (e.g., push-ups and pull-ups) and reinforcement learning (RL) policies (e.g., walking and turning), demonstrating the capability for zero-shot sim-to-real transfer. ToddlerBot can also acquire scalable real-world data. We design an intuitive teleoperation interface that allows simultaneous control of ToddlerBot’s upper and lower body to collect whole-body manipulation data and develop effective visuomotor policies. Additionally, ToddlerBot’s small size and weight ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensure safe and accessible operation in real-world environments.
Beyond ML-compatibility, ToddlerBot is designed with a focus on capability and reproducibility. A humanoid has the potential to utilize scalable human demonstrations by leveraging its anatomical similarity to the human body. As such, ToddlerBot features an anthropomorphic design with 30 active degrees of freedom (excluding end effectors), powered by carefully selected motors comparable to human muscle strengths normalized by the body size. We demonstrate ToddlerBot’s capabilities through arm span, payload, and endurance tests and various open-loop and closed-loop loco-manipulation tasks.
Reproducibility is achieved through low-cost, open-source designs and readily accessible hardware components. ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers). We will release digital twin software, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair. To validate reproducibility, we enlisted a CS-major student who is not involved in this project to independently build another instance of ToddlerBot with the provided assembly manual and successfully zero-shot transferred loco-manipulation policies between the two instances. Finally, we showcase a collaborative long-horizon task in which two ToddlerBot robots work together to tidy a room by organizing stuffed toys from a table and the floor.
II Related Works
@@
In recent years, numerous humanoid robots have been developed, showcasing diverse designs and capabilities. Industrial humanoids such as Boston Dynamics Atlas [8], Booster T1 [35], Cassie [33], Digit [34], Figure [9], Fourier GR1 [10], NAO H25 [31], Tesla Optimus [44], Robotis OP3 [36], Unitree G1 [47], and H1 [46] demonstrate remarkable technological advancements. On the other hand, humanoids from research institutions, including the Berkeley Humanoid [27], BRUCE [28], Duke Humanoid [49], iCub [32], and MIT Humanoid [5], also explore the humanoid design space with different emphasis. Humanoid robots can be evaluated using various metrics; we prioritize three core attributes: ML-compatibility, capability, and reproducibility. As outlined in Table I, we identify nine metrics to compare these attributes across different humanoid platforms.
@@
Size and weight are critical factors when designing humanoid robots. A smaller humanoid is inherently cheaper, easier to build and repair, and safer. Larger, full-size humanoids typically require a substantial engineering team for operation and maintenance, along with specialized facilities like gantry cranes for safety. In contrast, smaller humanoids can be deployed by a small team, often a single person, and operated in constrained environments with simply a laptop. But what is the trade-off with a smaller size? Reduced size can limit a robot’s ability to manipulate human-scale objects, However, miniature humanoids can still effectively perform meaningful manipulation tasks when paired with appropriately scaled objects. Furthermore, learning to manipulate smaller objects and developing dynamic whole-body control and locomotion techniques are likely transferable to larger humanoids.
@@
We argue that the number of active DoFs is crucial for the research value of a humanoid platform since more DoFs enable more human-like motion. While the human musculoskeletal system employs over 600 muscles working in complex synergies, the primary functional movements of the human body can be approximated using 32 revolute joints in a robotic system: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes. Therefore, humanoid designs aim to achieve a DoF count as close to 32 as possible. The passive DoFs are less important as they do not contribute to the action space. We believe that the public perceived limited performance of miniature humanoids is primarily due to fewer DoFs. This limitation often arises from space constraints that restrict the incorporation of many DoFs, a challenge we have successfully addressed in ToddlerBot.
@@
To qualitatively assess the capability, we evaluate the humanoid’s ability to perform both manipulation and locomotion tasks. While each type of motion is important, the combination of both is particularly compelling, as it unlocks opportunities for whole-body control research [19, 20, 12, 22, 29]. Furthermore, certain motions, such as push-ups, pull-ups, and cartwheels, go beyond traditional categories of manipulation and locomotion, treating them as prerequisites and requiring coordinated use of both arms and legs.
Recent advances show that large-scale simulation-based data collection is highly effective for locomotion [39, 42, 26], while real-world data collection is more promising for manipulation [6, 25, 30]. Therefore, an ideal humanoid research platform should facilitate data collection in both simulation and real-world settings, which we defined as ML-compatibility.
@@
Moreover, being open-source and low-cost is essential for others to reproduce. Without these qualities, research in this field would remain restricted to those with specialized expertise and significant resources. While making no compromise in functionality, ToddlerBot stands out as completely open-source and the most affordable among recent humanoid platforms, making it accessible to a wider range of researchers.
III System Design
Figure 2: Humanoid Metrics. Two key criteria of humanoid capability are the number of active DoFs and power factor $\tilde{p}$ (Equation 1). The total power factor is the sum of the upper and lower body power factor: $\tilde{p}_{\text{total}}=\tilde{p}_{\text{upper}}+\tilde{p}_{\text{lower}}$. ToddlerBot is the closest to human compared with other humanoids, implying potentially comparable loco-manipulation capabilities.
@@ III System Design
Figure 3: Mechatronic Design. We present three main sections: arm, leg, and head plus torso. We highlight ToddlerBot’s 30 active DoFs with orange markers: 7 DoFs per arm, 6 DoFs per leg, a 2-DoF neck, and a 2-DoF waist. Green markers indicate two end-effector designs—a compliant palm and a parallel-jaw gripper. Purple markers denote the sensor and electronics layout with exploded views, featuring two fisheye cameras, a speaker, two microphones, an IMU, and a Jetson Orin NX computer.
When designing humanoids, numerous factors must be considered, including payload, endurance, repeatability, degrees of freedom, cost, locomotion, and manipulation capabilities. Given this vast design space, ToddlerBot’s key design principles prioritize reproducibility, capability, and ML-compatibility. Unlike most prior works, we treat reproducibility as a hard constraint to narrow the design space, as our humanoid platform holds no value if it cannot be reproduced by others. With this assumption, we optimize the hardware for capability and enhance ML-compatibility through the careful design of the digital twin and teleoperation device.
III-A Reproducibility - A Hard Constraint
We define reproducibility as the ability of a single person to replicate the robot system at home without specialized equipment. Limited access to manpower and facilities presents several challenges, including fabrication, space, and cost.
Fabrication. 3D printing has become a popular method for reproducing open-source hardware systems due to its accessibility and fast turnaround time [24, 41, 48]. Despite these merits, a key challenge is to ensure that printed parts are strong enough for a humanoid robot. We find that PLA suits most parts, while high-strength components can be reinforced with a stronger printing profile and carbon fiber-blended filaments. To ensure ease of assembly and maintenance, we detail our design considerations in Section VIII-A of the supplementary material. For non-3D-printable components, such as motors, we limit ourselves to readily available off-the-shelf items.
Space. In a home environment without equipment to ensure safety, it is essential to minimize the robot’s size. However, the robot’s size is constrained by the onboard computer required for policy execution. We use the NVIDIA Jetson Orin NX for AI model execution, and to fit both the computer and battery within the torso while maintaining human-like proportions, the robot’s height is capped at approximately $0.6\leavevmode\nobreak\ \mathrm{m}$.
Cost. With 3D printing, the total BOM cost of ToddlerBot is $6000\leavevmode\nobreak\ \mathrm{USD}$, with 90% spent on the computers and motors.
III-B Capability - Design Objective
To quantitatively assess a humanoid robot’s capability, we propose two key metrics: power factor and the number of active DoFs. The power factor is defined as:
$\tilde{\mathit{p}}=\frac{\sum_{i=0}^{\mathit{N}}|{\tau}^{\max}_{i}|}{{h}\cdot{mg}},$ (1)
where $\mathit{N}$ denotes the number of active DoFs, $|{\tau}_{i}^{\max}|$ represents the peak torque of the $i$-th joint motor, $h$ denotes the humanoid’s height, and $mg$ represents the weight. We define $\tilde{\mathit{p}}$ as the power factor, representing the total torque (and thus mechanical power) a robot can generate relative to its weight and height scale. A detailed discussion of this metric is provided in Section VIII-B of the supplementary material.
Intuitively, having a higher $\tilde{\mathit{p}}$ means that a humanoid can perform energetic, dynamic motions more easily. We argue that $\tilde{\mathit{p}}$ should at least exceed the human threshold $\tilde{\mathit{p}}_{\text{human}}$ to achieve human-like motion, given the inherent gap between robot and human policies, assuming humans operate as an oracle policy that is energy efficient. However, raising $\tilde{\mathit{p}}$ far beyond $\tilde{\mathit{p}}_{\text{human}}$ can have adverse effects: it may lead to unnatural motion, excessive reliance on motor power, fewer DoFs to accommodate larger motors, reduced battery life, and increased safety concerns. Thus, pushing $\tilde{\mathit{p}}$ past diminishing returns involves a practical trade-off. As shown in Figure 2, ToddlerBot has a $\tilde{\mathit{p}}$ score closest to humans.

### v3

@@ Abstract
Learning-based robotics research driven by data demands a new approach to robot hardware design—one that serves as both a platform for policy execution and a tool for embodied data collection. We introduce ToddlerBot, a low-cost, open-source humanoid robot platform designed for robotics and AI research. ToddlerBot enables seamless acquisition of high-quality simulation and real-world data. The plug-and-play zero-point calibration and transferable motor system identification ensure a high-fidelity digital twin and zero-shot sim-to-real policy transfer. A user-friendly teleoperation interface streamlines real-world data collection from human demonstrations. With its data collection ability and anthropomorphic design, ToddlerBot is ideal for whole-body loco-manipulation research. Additionally, ToddlerBot’s compact size ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensures safe operation in real-world environments. Reproducibility is achieved with entirely 3D-printed, open-source design and off-the-shelf components, keeping the total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$. This allows assembly and maintenance with basic technical expertise, as validated by successful independent replications of the system. We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, loco-manipulation tasks, and a collaborative long-horizon scenario where two robots tidy a toy session together. By advancing ML-compatibility, capability, and reproducibility, ToddlerBot provides a robust and scalable platform for policy learning and execution in robotics research.
Keywords: Humanoid, Mechanisms & Design, Robot Modeling & Simulation
@@
Figure 1: ToddlerBot is an open-source humanoid platform for large-scale, high-quality data collection in both simulation and the real world. It combines massive parallel simulation and an accurate digital twin for simulation, and an intuitive teleoperation device for whole-body control in the real world. ToddlerBot enables diverse loco-manipulation skills, including walking, push-ups, pull-ups, wagon pushing, bimanual, and full-body manipulation, learned from both data sources.
1 Introduction
Conventional robot design prioritizes actuator strength, sensor accuracy, and mechanical precision. While these characteristics are essential for deploying robots in real-world environments, they fall short in addressing the needs of robotics research—where affordability, rapid repairability, and full-stack ownership with no black boxes are crucial for faster innovation, easier debugging, and total system understanding. Additionally, traditional robot platforms are often not aligned with modern machine learning (ML) paradigms driven by embodied data. A robot platform compatible with ML must possess the innate ability to collect observation and action data, both in simulation and the real world, as these complementary data sources are essential for scalable policy learning.
In response to these challenges, we introduce ToddlerBot, an open-source humanoid robot platform for the robotics and AI research community (Figure 1), specifically developed to facilitate policy learning for locomotion and manipulation skills. ToddlerBot is designed to resemble a miniature human to minimize construction and maintenance costs and specialized expertise, lowering the barrier for researchers seeking to build self-sufficient, fully transparent humanoids free from critical reliance on third-party components. Although ToddlerBot lacks the payload capacity of full-sized humanoids, it offers more degrees of freedom in both the upper and lower body than any existing miniature humanoids—sufficient for advancing research in manipulation and locomotion.
Inspired by recent quadrupeds [1, 2] and robotic manipulators [3, 4, 5, 6, 7], ToddlerBot is designed with ML-compatibility in mind. Researchers can use ToddlerBot to not only execute policies but also to serve as a robust data collection platform. It enables the acquisition of high-quality simulation data through a plug-and-play zero-point calibration procedure and transferable motor system identification (sysID) results. These tools ensure a high-fidelity digital twin. We validate the quality of simulation data with keyframe-interpolated motions (e.g., push-ups and pull-ups) and reinforcement learning (RL) policies (e.g., walking), demonstrating the capability for zero-shot sim-to-real transfer. ToddlerBot can also acquire scalable real-world data. We design an intuitive teleoperation interface that allows simultaneous control of ToddlerBot’s upper and lower body to collect whole-body manipulation data and develop effective visuomotor policies. Additionally, ToddlerBot’s small size and weight ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensure safe and accessible operation in real-world environments.
Beyond ML-compatibility, ToddlerBot is designed with a focus on capability and reproducibility. A humanoid has the potential to utilize large scale human demonstrations by leveraging its anatomical similarity to the human body. As such, ToddlerBot features an anthropomorphic design with 30 active degrees of freedom (excluding end effectors). We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, and various open-loop and closed-loop loco-manipulation tasks.
Reproducibility is achieved through low-cost, open-source designs and readily accessible hardware components. ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers). We release digital twin, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair. To validate reproducibility, we enlisted a CS-major student not involved in this project to independently build another instance of ToddlerBot. and successfully zero-shot transfer loco-manipulation policies between the two instances. We also received five reports of independent replications of ToddlerBot within a week by teams worldwide.
We summarize our contributions as follows: (1) As a system paper, we integrate existing components to create the first miniature humanoid with 30 DoFs and a superhuman range of motion, offering a practical humanoid platform with new capabilities for the community. (2) We develop a comprehensive sysID pipeline to ensure a high-fidelity digital twin in simulation. (3) We design a whole-body teleoperation solution to collect loco-manipulation data in the real world.
2 Related Works
Table 1: Comparison with Other Popular Humanoid Research Platforms.
@@ 2 Related Works
Humanoid Size Weight Compute Active DoFs(b) Manipulation Locomotion Sim Data Real Data Open Source Price
(m) (kg) (TFLOPS)(a) - - - - - - (\$)
BD Atlas [8] 1.50 89.0 - 28 - - -
Figure [9] 1.68 70.0 - 26 - - -
Optimus [10] 1.73 57.0 - 28 - - -
Digit [11] 1.75 65.0 - 16 code 250K
Unitree H1 [12] 1.76 47.0 1.92 19 code 70K
Fourier GR1 [13] 1.65 55.0 2.23 32 - - 110K
Booster T1 [14] 1.18 30.0 3.33 23 code 34K
Unitree G1 [15] 1.32 35.0 2.50 29 code 57K
iCub [16] 1.04 24.0 1.93 32 code 300K
Berkeley [17] 0.85 16.0 1.92 12 code 10K
MIT [18] 1.04 24.0 - 18 - - -
Duke [19] 1.00 30.0 Optional 10 design, code 16K
Berkeley Lite [20] 0.80 16.0 0.29 22 design, code 5K
BRUCE [21] 0.70 4.8 0.1 16 code 6.5K
NAO H25 [22] 0.57 5.2 0.02 23 code 14K
Robotis OP3 [23] 0.51 3.5 0.30 20 code 11K
Zeroth [24] 0.48 3.6 0.01 16 design, code 1.4K
Ours 0.56 3.4 2.50 30 design, code 6K
Average Adult [25] 1.73 70.9 - 32(c) - - - -
@@
    TFLOPS refers to Tera FLoating-point Operations Per Second with single precision (FP32).
  • (b)
@@
    While human body is powered by over 600 muscles, the primary functional movements can be approximated using 32 revolute joints: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes.
@@
In recent years, numerous humanoid robots have been developed. Full size industrial humanoids such as Boston Dynamics Atlas [8], Cassie [26], Digit [11], Figure [9], Tesla Optimus [10], Unitree H1 [12] and Fourier GR1 [13]. Half-scaled Booster T1 [14], Berkeley Humanoid Lite  [20] and Unitree G1 [15]. Miniature-sized NAO H25 [22], Robotis OP3 [23], and K-Scale Zeroth [24]. Together, they showcase diverse designs and capabilities. On the other hand, humanoids from research institutions, including the Berkeley Humanoid [17], BRUCE [21], Duke Humanoid [19], iCub [16], and MIT Humanoid [18], also explore the humanoid design space with different optimization emphasis. As outlined in Table 1, we identify nine metrics to compare different humanoid platforms.
@@
Size and weight are critical when designing humanoid robots. For manipulation, full-size humanoids are compatible with human-scale objects, whereas reduced-size robots are limited in their ability to manipulate them. However, reduced-size humanoids can still benefit manipulation research when paired with appropriately scaled objects [20]. Furthermore, dynamic whole-body control techniques are likely transferable to larger humanoids. For locomotion, full-sized and small-scale humanoids offer similar research values [27]. Full-size humanoids typically require a substantial engineering team to operate and maintain, along with specialized facilities like gantry cranes for safety. In contrast, a smaller humanoid is inherently cheaper, easier to build and repair, and safer [28, 27, 29]. They can be deployed by a single person and operated in constrained environments with a laptop.
@@
The number of active DoFs is crucial for human-like motion. While the human musculoskeletal system employs over 600 muscles, primary functional movement can be approximated using 32 revolute joints [16, 30]: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes. Therefore, humanoid designs aim to achieve a DoF count as close to 32 as possible. The public perception of the limited performance of miniature humanoids is primarily due to fewer DoFs. This limitation often arises from space constraints that restrict the incorporation of many DoFs, a challenge we have successfully addressed in ToddlerBot.
@@
To qualitatively assess the capability, we evaluate the humanoid’s ability to perform both manipulation and locomotion tasks. The combination of both is particularly compelling, as it unlocks opportunities for whole-body control research [31, 32, 33, 34, 35]. Furthermore, certain motions, such as push-ups, pull-ups, and cartwheels, require coordinated use of both arms and legs. Apart from physical capabilities, ToddlerBot also has best compute in class. Onboard CUDA accelerator allows concurrent policy inferences to support visual and locomotion capabilities. Large-scale simulation data collection is effective for locomotion [36, 37, 38], while real-world data collection is more promising for manipulation [39, 40, 41]. Therefore, to be ML-compatible, a humanoid research platform should facilitate data collection in both world.
@@
Moreover, being open-source and low-cost is essential for others to reproduce. Without these qualities, research in this field would remain restricted to those with specialized expertise and significant resources. While making no compromise in functionality, ToddlerBot stands out as completely open-source and the most accessible among recent humanoid research platforms.
3 System Design
The design space of humanoid robots is vast; ToddlerBot’s key design principles prioritize reproducibility as a hard constraint, and capability and ML-compatibility as key design objectives.
3.1 Reproducibility - A Hard Constraint
Unlike most prior works, we treat reproducibility as a hard constraint, as our humanoid platform holds no value if others cannot reproduce it. We define reproducibility as the ability of a single person to replicate the robot system at home without specialized equipment. 3D printing has become a popular method for reproducing open-source hardware systems due to its accessibility and fast turnaround time [2, 5, 4]. Despite these merits, it’s challenging to ensure sufficient strength and precision for printed plastic parts. We discuss our justifications for 3D printing in Appendix 8.1. For performance-critical components like motors and bearings, we limit ourselves to off-the-shelf parts. The total BOM cost of ToddlerBot is $6,000\leavevmode\nobreak\ \mathrm{USD}$, with 90% spent on the computer (NVIDIA Jetson) and motors (Robotis Dynamixel). We detail our design considerations in Appendix 8.2.
@@ 3.1 Reproducibility - A Hard Constraint
Figure 2: Mechatronic Design. Orange markers highlights ToddlerBot’s 30 active DoFs: 7 per arm, 6 per leg, a 2 on neck, and a 2 on waist. Green markers indicate two end-effector designs—a compliant palm and a parallel-jaw gripper. Purple markers denote the electronics layout with exploded views, featuring 2 fisheye cameras, 1 speaker, 2 microphones, 1 IMU, and 1 Jetson Orin computer.
3.2 Capability - Design Objective